Circular No. 7249

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CASSIOPEIA A B. Aschenbach, Max-Planck-Institut für Extraterrestrische Physik,

Garching, reports: “The analysis of a 180 000-s observation of Cas A taken with the ROSAT HRI in 1995–96 shows a pointlike source at α = 23^{h}23^{m}27^{s}.6, δ = +58^{o}48′44′′.0 (equinox 2000.0) that appears to be the same object reported by Tananbaum et al. from the Chandra First Light observation (IAUC 7246), because of the small separation of the two sources (< 2′′) and the absence of other candidates. The ROSAT HRI backgroundcorrected countrate was 0.0015 ± 0.0003 cts/s. Correction for interstellar absorption is highly uncertain because of the large N_{H}. However, taking “standard” values of N_{H} = 1.2 × 10^{22} cm^{−2} and a distance of 3.4 kpc, the countrate is consistent with blackbody emission from a 10-km radius neutron star with a temperature of 1.6 MK, and this supports the view that the object is the Cas A neutron star.”

NOVA IN M31 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126), reports the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a new nova, of mag ∼ 17.0, in an unfiltered image of M31 taken on Sept. 2.3 UT. The nova was confirmed on earlier images taken on Aug. 30.3, 31.3 and Sept. 1.3 at mag ∼ 18.5, 17.8 and 17.0, respectively. It is located at α = 0^{h}42^{m}41^{s}.1, δ = +41^{o}19′11′′ (equinox 2000.0), which is ∼ 43′′ west and 175′′ north of the nucleus of M31. A KAIT image on Aug. 28.3 showed nothing at the position of the nova (limiting mag 19.0). A finding chart

can be found at http://astron.berkeley.edu/~bait/1999/n31 3.html.

NOVA CIRCINI 1999 Photometry by P. M. Kilmartin and A. C. Gilmore with the 0.6-m f/16

reflector at Mt. John Observatory: Aug. 26.360 UT, V = 8.49, U − B = −0.69, B − V = +0.53, V − R = +0.90, V − I = +1.58, airmass = 1.28; 27.387, 9.58, −0.63, +0.47, +1.32, +1.81, 1.35; 31.359, 9.74, −0.50, +0.35, +1.63, +1.68, 1.31. Reference stars for Aug. 26 were Cousins’ F213 = HD 107547 and F210 = HD 106759. Nearby reference stars for Aug. 27 and 31, with assumed magnitudes and colors, calibrated from Cousins’ F-region standards, were HD 125389 (V = 7.35, U − B = −0.07, B − V = +0.43, V −R = +0.25, V −I = +0.52) and HD 125605 (9.09, +0.04, +0.56, +0.33, +0.68). Standard deviations are 0.01 mag or less.

1999 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden